# A review of regularised estimation methods and cross-validation in spatiotemporal statistics

### v1

@@ 3.2 Regularisation for geospatial interactions
Second, suppose that the random effects follow a $q$-variate Gaussian process, i.e.,
@@
where $C_{\theta}$ is a covariance function that depends on the difference between any two locations $\boldsymbol{s}$ and $\boldsymbol{s}^{\prime}$ and two arbitrary time points $t$ and $t^{\prime}$. Moreover, if $q>1$, $C_{\theta}$ gives a $q\times q$ covariance matrix for cross-variable correlation. This covariance function can depend on further unknown parameters $\theta$, which have to be estimated. Here, in this case, the covariance function is called separable (Cressie and Wikle,, 2015). Moreover, if $C_{\theta}$ is only a function of $||\boldsymbol{s}-\boldsymbol{s}^{\prime}||$ with |||||$\cdot$||| | ⋅ | | being a vector norm (i.e., if only the distance between $\boldsymbol{s}$ and $\boldsymbol{s}^{\prime}$ matters, but their orientation towards each other), the covariance function is isotropic. Generally, it is desirable if the covariance matrix resulting from $C_{\theta}$ contains many zeros; that is, it is sparse. The traditional way to induce zeros in the covariance is known as covariance tapering (Furrer et al., 2006b, ). Loosely speaking, based on the geographical distance between the locations, a covariance of zero is assumed for observations whose distance is larger than a certain threshold (theoretical results on covariance tapering can be found in Stein, 2013). This method can also be applied in likelihood-based estimation procedures (see Kaufman et al., 2008, and Furrer et al., 2016 for theoretical results). As mentioned above, regularised estimation procedures are tailor-made to induce zeros for certain parameters when the shrinkage target is chosen to be zero. Typically, we can find zeros if two observations across space/time are conditionally independent (i.e., independent when observing all other realisations). However, the conditional independence is encoded in the inverse covariance matrix or precision matrix. Thus, penalised methods can be applied to obtain sparse precision matrices. For univariate data ($q=1$), Krock et al., 2021a introduced a graphical LASSO procedure to induce zeros in the precision matrix of a spatial process. Krock et al., 2021b extended this approach to be used in the multivariate case ($q>1$).
@@
Moreover, since the covariance matrix is a positive definite matrix by definition (i.e., for valid choices $C_{\theta}$, see Gneiting, 2002; Gneiting et al., 2010; Stein, 2005; Nychka et al., 2002; Porcu et al., 2016 or Porcu et al., 2021 for a historical review of space-time covariance functions), both the covariance matrix and its inverse can be decomposed as $\mathbf{\Sigma}=\mathbf{P}\mathbf{P}^{\prime}$, e.g., via Cholesky decomposition. The matrix $\mathbf{P}$ is called Cholesky factor. Stein et al., 2004b proposed to approximate the likelihood for large spatial data sets based on Vecchia approximations (Vecchia,, 1988). The idea is to approximate the joint likelihood as a product of the conditional likelihoods. Now, sparse Cholesky factors have to be considered first by Schäfer et al., (2021) for spatial models. If the dimension of $\mathbf{P}$ is $N\times r$ with $N$ being the total number of observations across space/time and $r\ll N$, the covariance matrix is low-rank. That is, the dependence structure reduces to a lower-dimensional space. In spatial and spatiotemporal statistics, low-rank covariance matrices have been considered first by Banerjee et al., (2008), Cressie and Johannesson, (2008) (spatial fixed-rank kriging), and Cressie et al., (2010) (spatiotemporal fixed-rank filtering). Back to the subject of this review paper, Chang et al., (2010) proposed a penalised estimation procedure to identify the lower rank of the covariance matrix. Specifically, low-rank approximations aim to represent a spatiotemporal process as a linear combination of local basis functions, which are weighted by uncorrelated random-effect coefficients (see also the recent review by Cressie et al., 2022). Thus, regularised estimation procedures aim to select suitable local basis functions. From a practical perspective, many different local basis functions (e.g., on several grids with different resolutions) can be included, and the best basis functions are chosen automatically by estimating the model parameters. For the spatiotemporal case, Hsu et al., (2012) suggested penalised procedures to choose these local basis functions. Furthermore, Kang and Katzfuss, (2021) considered sparse inverse Cholesky factors that are identified based on correlations.
@@ 3.3 Bayesian estimation procedures
Along with the frequentist paradigm, the literature pioneered penalised regression extensions following a Bayesian perspective. Bayesian estimation schemes have gained particular importance for spatiotemporal models, especially using integrated nested Laplace approximations (INLA) which makes them applicable also for large data sets (see Rue et al.,, 2017, for a review on INLA in the spatiotemporal context). Moreover, we refer the readers to the review paper by van Erp et al., (2019) for a comprehensive overview of the state-of-the-art literature on Bayesian penalised regression. The authors summarise that Bayesian penalisation techniques include the penalised in three alternative ways. The first way is called Full Bayes or hierarchical Bayes (Wolpert and Strauss,, 1996) approach, which treats the penalty parameter $\lambda$ as an unknown variable (i.e., a hyperparameter) whose prior distribution has to be specified. Such models specify prior distributions for all parameters and can be estimated in a single step. The prior distribution is called the shrinkage prior (van Erp et al.,, 2019) and is usually a vague distribution, e.g., a half-Cauchy random variable (Polson and Scott,, 2012). The prior acts on the coefficients in order to shrink small effects to zero while maintaining true large effects. Indeed, large values of $\lambda$ result in smaller prior variation and thus more shrinkage of the coefficients towards zero.
The second way is named Empirical Bayes (van de Wiel et al.,, 2019) and intends the parameters as unknown constants. This approach differs from the first one because it involves a two-step process: first, estimating the penalty parameter $\lambda$ from the observed data, and second, incorporating this empirical estimate into the model using an empirical Bayes prior distribution. Since the empirical approach does not specify any prior distribution for the hyperparameters, a sensitivity analysis of the results with respect to the distributions is not necessary.
@@
In addition to the penalised regression approach discussed above, the Bayesian framework includes other model selection techniques, such as the Zellner’s g-prior (Zellner,, 1986), and the spike-and-slab prior (Mitchell and Beauchamp,, 1988). The former approach shrinks the regression coefficients toward zero through a global shrinkage scalar called g, which equally shrinks each coefficient (which can be reasonable if the coefficients are equivalent). The single-g Zellner’s prior has been extended in several ways, e.g. using a mixture of g priors (Liang et al.,, 2008), and multiple shrinkage factors as in Zhang et al., (2016). However, carefully choosing the constant g must be addressed to avoid excluding important variables (Lindley,, 1957). Instead, according to the data, the spike-and-slab method assigns the regression coefficients to the zero-centred spike (i.e., shrinking toward zero) if they do not deviate substantially from zero. In contrast, if they differ significantly from zero, they will be assigned to the slab (i.e., the vague proper prior). Both methods were also adapted to spatio-temporal analysis. Refer, for example, to Lee et al., (2014) on the combination of Zellner’s g prior and spatial Ising prior for selecting spatial covariates in spatial time series data.
@@
Possible alternatives to penalty methods for selecting linear predictors in space-time models are mixture model selection methods. This category includes Bayesian selection methods and Bayesian model averaging. The former evaluates the appropriateness of a model based on the estimated weight among a variety of models with alternative predictors; the latter, on the other hand, averages over several alternative models to find the posterior distribution of the parameters. Spatial dynamics is included in the algorithms by adding an intrinsic conditional autoregressive (ICAR) (Besag et al.,, 1991; Besag and Green,, 1993) approach (Carroll et al.,, 2018), while temporal dynamics is incorporated via autoregressive processes (Lawson et al.,, 2017). Both methods have proven effective in disease mapping studies with spatial data (Carroll et al.,, 2018), spatial small area frameworks (Carroll et al., 2016a, ), as well as spatio-temporal disease mapping (Carroll et al., 2016b, ). Among others, relevant advantages include getting an automatic final model fit.
@@ 4 Spatial autoregression
$\boldsymbol{Y}_{t}=\mathbf{X}_{t}\boldsymbol{\beta}+\rho\mathbf{W}\boldsymbol{Y}_{t}+\boldsymbol{\varepsilon}_{t}\,,$ (3)
where $\mathbf{X}_{t}$ is a matrix of regressors, $\boldsymbol{\beta}$ is the corresponding vector of regression coefficients, $\rho$ is the spatial autoregressive parameter, and $\boldsymbol{\varepsilon}_{t}$ is the vector of model errors. For spatial autoregressive (SAR) and conditional autoregressive (CAR) models, Gonella et al., (2022) proposed a LASSO estimation procedure to select the relevant covariates in the regression term.
@@
Alternatively, the spatial interactions can be modelled in the random effects term in the same manner, while the errors remain independent across space. Generally, there is a relation between geostatistical and spatial autoregression models, and both approaches are equivalent under certain conditions (Ver Hoef et al.,, 2018, Theorem 1).
@@
Instead of the linear relation $\rho\mathbf{W}$, the spatial interactions can be modelled using a series of different weight structures, e.g., $\sum_{i=1}^{k}\rho_{i}\mathbf{W}_{i}$ with $k$ different weight matrices $\mathbf{W}_{i}$. For instance, each weight matrix could only contain the weights for certain directions (northward, north-eastward, eastward dependence, etc.) to reveal directional processes (Merk and Otto,, 2021). Moreover, penalised estimation procedures can be used to select the true weight matrix $\mathbf{W}$ from a series of alternative weights $\mathbf{W}_{1},\ldots,\mathbf{W}_{k}$, as for the boosting procedure proposed by Kostov, (2010, 2013) or the LASSO least-squares procedure proposed by Lam and Souza, (2020). Reyes et al., 2012b applied a spatiotemporal LASSO procedure to select weight matrices from a set of candidates with increasing spatial lag order, simultaneously with the temporal lags and spatiotemporal weight matrices. In other words, they constructed a 2-dimensional grid of the temporal and spatial lag orders and selected the relevant spatiotemporal interactions. In the context of traffic analysis, Haworth and Cheng, (2014) applied a graphical LASSO procedure to select local neighbourhood structures analogously. Together with the autoregressive coefficients, Reyes et al., 2012b also penalised the regressive parameters using a second penalty term. Similarly, Liu et al., (2018) suggested a LASSO procedure for selecting the regressors in the mean equation while allowing for spatial autoregressive structure in the model. Since ordinary least-squares procedures are inconsistent in the presence of spatial autoregressive dependence, they proposed a penalised quasi-maximum likelihood approach incorporating a LASSO penalty with a zero shrinkage target for the regression coefficients.
@@ 5 Cross validation under spatiotemporal dependence
When a strictly positive $\lambda$ is used, the estimated coefficients $\beta_{\lambda}$ are biased but are more efficient, i.e., their variance is smaller. This is known as the bias-variance tradeoff. Additionally, opting for a $\lambda$ value that is too small can result in overfitting for richly parametrised models, while selecting a value that is too large can lead to underfitting (see, e.g., Boonstra et al.,, 2015). The optimal penalty parameter is usually selected based on the model’s in-sample or out-of-sample predictive performance. Alternative ways to select the penalty parameter are likelihood-based methods, where $\lambda$ is interpreted as a variance component and the likelihood is maximised with respect to the couple $(\sigma_{\varepsilon}^{2},\lambda)^{\prime}$. Similarly, the penalty parameter can be simultaneously estimated with all parameters in a fully Bayesian approach, where typically vague half-Cauchy prior distribution is assumed for $\lambda$ (van Erp et al.,, 2019). Alternatively, Otto and Steinert, (2022) proposed to select the penalty parameter based on the distance between the sample and model spatial autocorrelation. Below, we will focus on the predominantly applied goodness-of-fit (GoF) criteria.
GoF criteria aim to maximise the model fit by tuning the penalty parameter $\lambda$. The model fit could be assessed by the in-sample fit in terms of distance between the observed and predicted values or the out-of-sample fit, i.e., cross-validation prediction accuracies. The distance is typically evaluated by Root-Mean-Squared-Error (RMSE), Mean Absolute Error (MAE), or based on information criteria, such as Akaike’s Information Criterion (Akaike,, 1973, AIC) and the Bayesian Information Criterion (Schwarz,, 1978, BIC). Most models have no closed-form solutions to determine the optimal value of $\lambda$, necessitating grid-search algorithms. Arlot and Celisse, (2010) offers a thorough review of the significance of cross-validation (CV) in regression model selection and its statistical properties.
@@
    mutual independence of test errors
@@
    independence between the training set and test set.
@@
Thus, depending on the data structure (e.g., cross-sectional, clustered, spatial, temporal data), the cross-validation schemes to be adopted may vary significantly. In the case of independent data, classical CV schemes, such as random $k$-fold, stratified $k$-fold (Zeng and Martinez,, 2000; Ludwig et al.,, 2016) or Generalized CV (Boonstra et al.,, 2015) can be used. For instance, considering ridge regularisation, the GCV estimator can be used to efficiently estimate the penalty parameter $\lambda$ even when the number of observations (thus, the degrees of freedom) is small or the number of parameters to be estimated exceeds the number of observations (Golub et al.,, 1979).
In the case of spatially/temporally dependent data, standard random $k$-fold cross-validation procedures cannot be applied. Schratz et al., (2019) showed that, while performances are overestimated when no spatial information is included in the CV step, hyperparameter tuning of machine learning models appears to be less sensitive to the spatial structure, leading to similar results of non-spatial and spatial CV schemes. The problems arising from the occurrence of spatio-temporal autocorrelation in model assessment are manifold.
@@
First, using random sampling in cross-validation results in test observations being collected from areas that are spatially close to the training observations (Schratz et al.,, 2019). Consequently, the evaluation of prediction performance tends to be overly optimistic because the training and test datasets become correlated, largely due to the neglected underlying correlation structure, whether across space or time (Brenning,, 2012; Meyer et al.,, 2019; Meyer and Pebesma,, 2021; Ploton et al.,, 2020; Lezama Valdes et al.,, 2021). That is, if one leaves out a set of observations and estimates a spatiotemporal model with the remaining observations, information from the observations used for model estimation is used to predict the left-out observations (via the spatiotemporal interactions). Second, when the spatiotemporal structure is neglected, the estimated residuals will not be mutually independent, which is a critical assumption in many statistical models. As a result, it is often advisable to exclude complete data blocks across time and/or space (Roberts et al.,, 2017).

### v2

@@ 3.2 Regularisation for the random-effects component
Suppose that the random-effects component follow a stationary $q$-variate Gaussian process, i.e.,
@@
where $C_{\theta}$ is a matrix covariance function that depends on the difference between any two locations $\boldsymbol{s}$ and $\boldsymbol{s}^{\prime}$ and two arbitrary time points $t$ and $t^{\prime}$. It is the fundamental building block of the covariance matrix of the data set (1) and, for $q>1$, also provides the cross-covariances between the components of the observation vector $y$. In order to have a positive semidefinite observation covariance matrix, $C_{\theta}$ must be a “valid” covariance function. See Gneiting, 2002; Gneiting et al., 2010; Stein, 2005; Nychka et al., 2002; Porcu et al., 2016 or Porcu et al., 2021 for a historical review of space-time covariance functions. The covariance function in (3) usually depends on some unknown parameters $\theta$, which have to be estimated.
@@
Generally, it is desirable if the covariance matrix resulting from $C_{\theta}$ contains many zeros; that is, it is sparse. The traditional way to induce zeros in the covariance is known as covariance tapering (Furrer et al., 2006b, ). Loosely speaking, based on the geographical distance between the locations, a covariance of zero is assumed for observations whose distance is larger than a certain threshold. Theoretical results on covariance tapering can be found in Stein, 2013, and the use of tapering in the multivariate case is considered in Bevilacqua et al., (2016). This method can also be applied in likelihood-based estimation procedures (see Kaufman et al., 2008, and Furrer et al., 2016 for theoretical results). As mentioned above, regularised estimation procedures are tailor-made to induce zeros for certain parameters when the shrinkage target is chosen to be zero. Typically, we can find zeros if two observations across space/time are conditionally independent (i.e., independent when observing all other realisations). However, the conditional independence is encoded in the inverse covariance matrix or precision matrix. Thus, penalised methods can be applied to obtain sparse precision matrices. For univariate data ($q=1$), Krock et al., 2021a introduced a graphical LASSO procedure to induce zeros in the precision matrix of a spatial process. Krock et al., 2021b extended this approach to be used in the multivariate case ($q>1$).
@@
Moreover, since the covariance matrix is a positive definite matrix by definition, both the covariance matrix and its inverse can be decomposed as $\mathbf{\Sigma}=\mathbf{P}\mathbf{P}^{\prime}$, e.g., via Cholesky decomposition. The matrix $\mathbf{P}$ is called Cholesky factor. Stein et al., 2004b proposed to approximate the likelihood for large spatial data sets based on Vecchia approximations (Vecchia,, 1988). The idea is to approximate the joint likelihood as a product of the conditional likelihoods. Now, sparse Cholesky factors have to be considered first by Schäfer et al., (2021) for spatial models. If the dimension of $\mathbf{P}$ is $N\times r$ with $N$ being the total number of observations across space/time and $r\ll N$, the covariance matrix is low-rank. That is, the dependence structure reduces to a lower-dimensional space. In spatial and spatiotemporal statistics, low-rank covariance matrices have been considered first by Banerjee et al., (2008), Cressie and Johannesson, (2008) (spatial fixed-rank kriging), and Cressie et al., (2010) (spatiotemporal fixed-rank filtering). Back to the subject of this review paper, Chang et al., (2010) proposed a penalised estimation procedure to identify the lower rank of the covariance matrix. Specifically, low-rank approximations aim to represent a spatiotemporal process as a linear combination of local basis functions, which are weighted by uncorrelated random-effect coefficients (see also the recent review by Cressie et al., 2022). Thus, regularised estimation procedures aim to select suitable local basis functions. From a practical perspective, many different local basis functions (e.g., on several grids with different resolutions) can be included, and the best basis functions are chosen automatically by estimating the model parameters. For the spatiotemporal case, Hsu et al., (2012) suggested penalised procedures to choose these local basis functions. Furthermore, Kang and Katzfuss, (2021) considered sparse inverse Cholesky factors that are identified based on correlations.
@@ 3.3 Bayesian estimation procedures
Along with the frequentist paradigm, the literature pioneered penalised regression extensions following a Bayesian perspective. Bayesian estimation schemes have gained particular importance for spatiotemporal models, especially using integrated nested Laplace approximations (INLA), which makes them applicable also for large data sets (see Rue et al.,, 2017, for a review on INLA in the spatiotemporal context). Moreover, we refer the readers to the review paper by van Erp et al., (2019) for a comprehensive overview of the state-of-the-art literature on Bayesian penalised regression. The authors summarise that Bayesian penalisation techniques include the penalised in three alternative ways. The first way is called full Bayesian or hierarchical Bayesian (Wolpert and Strauss,, 1996) approach, which treats the penalty parameter $\lambda$ as an unknown variable (i.e., a hyperparameter) whose prior distribution has to be specified. Such models specify prior distributions for all parameters and can be estimated in a single step. The prior distribution is called the shrinkage prior (van Erp et al.,, 2019) and is usually a vague distribution, e.g., a half-Cauchy random variable (Polson and Scott,, 2012). The prior acts on the coefficients in order to shrink small effects to zero while maintaining true large effects. Indeed, large values of $\lambda$ result in smaller prior variation and thus more shrinkage of the coefficients towards zero.
The second way is named empirical Bayesian approach (van de Wiel et al.,, 2019) and intends the parameters as unknown constants. This approach differs from the first one because it involves a two-step process: first, estimating the penalty parameter $\lambda$ from the observed data, and second, incorporating this empirical estimate into the model using an empirical Bayes prior distribution. Since the empirical approach does not specify any prior distribution for the hyperparameters, a sensitivity analysis of the results with respect to the distributions is not necessary.
@@
In addition to the penalised regression approach discussed above, the Bayesian framework includes other model selection techniques, such as the Zellner’s g-prior (Zellner,, 1986), and the spike-and-slab prior (Mitchell and Beauchamp,, 1988). The former approach shrinks the regression coefficients toward zero through a global shrinkage scalar called g, which equally shrinks each coefficient (which can be reasonable if the coefficients are equivalent). The single-g Zellner’s prior has been extended in several ways, e.g. using a mixture of g priors (Liang et al.,, 2008), and multiple shrinkage factors as in Zhang et al., (2016). However, carefully choosing the constant g must be addressed to avoid excluding important variables (Lindley,, 1957). Instead, according to the data, the spike-and-slab method assigns the regression coefficients to the zero-centred spike (i.e., shrinking toward zero) if they do not deviate substantially from zero. In contrast, if they differ significantly from zero, they will be assigned to the slab (i.e., the vague proper prior). Both methods were also adapted to spatiotemporal analysis. Refer, for example, to Lee et al., (2014) on the combination of Zellner’s g prior and spatial Ising prior for selecting spatial covariates in spatial time series data.
@@
Possible alternatives to penalty methods for selecting linear predictors in space-time models are mixture model selection methods. This category includes Bayesian selection methods and Bayesian model averaging. The former evaluates the appropriateness of a model based on the estimated weight among a variety of models with alternative predictors; the latter, on the other hand, averages over several alternative models to find the posterior distribution of the parameters. Spatial dynamics is included in the algorithms by adding an intrinsic conditional autoregressive (ICAR) (Besag et al.,, 1991; Besag and Green,, 1993) approach (Carroll et al.,, 2018), while temporal dynamics is incorporated via autoregressive processes (Lawson et al.,, 2017). Both methods have proven effective in disease mapping studies with spatial data (Carroll et al.,, 2018), spatial small area frameworks (Carroll et al., 2016a, ), as well as spatiotemporal disease mapping (Carroll et al., 2016b, ). Among others, relevant advantages include getting an automatic final model fit.
@@ 4 Spatial autoregression
$\boldsymbol{Y}_{t}=\mathbf{X}_{t}\boldsymbol{\beta}+\rho\mathbf{W}\boldsymbol{Y}_{t}+\boldsymbol{\varepsilon}_{t}\qquad\text{for $t=1,\ldots,T$}\,,$ (4)
where $\mathbf{X}_{t}$ is a matrix of regressors, $\boldsymbol{\beta}$ is the corresponding vector of regression coefficients, $\rho$ is the spatial autoregressive parameter, and $\boldsymbol{\varepsilon}_{t}$ is the vector of white noise model errors.
4.1 Regularisation for the mean model
For spatial autoregressive (SAR) and conditional autoregressive (CAR) models, Gonella et al., (2022) proposed a LASSO estimation procedure to select the relevant covariates in the regression term. Among others, Wen et al., (2018) use a spatial autoregressive model with adaptive LASSO penalty to detect relevant autoregressive parameters in genetic studies. Zhu et al., (2010) implemented an iterative penalised likelihood estimator with adaptive LASSO penalty to select predictors and neighbourhood structure in conditional autoregressive and simultaneous autoregressive models with spatially correlated error terms. Later on, Reyes et al., (2012) proposed an adaptive LASSO algorithm for the case of linear regression models with spatiotemporal neighbourhood structures. Liu, (2017) extended the previous algorithms, allowing for spatial correlation to be captured by either the spatial lag terms or spatial errors or both through a SARAR model. Also, their penalised estimates are obtained via least squares approximation to account for possible non-concavity of the likelihood function. Other examples of penalised likelihood for spatiotemporal data are in Al-Sulami et al., (2019), in which an adaptive LASSO method is proposed to simultaneously identify and estimate spatiotemporal lag interactions in the context of a data-driven semiparametric nonlinear model. Similarly, Liu, (2022) developed an adaptive LASSO variable selection method for semiparametric spatial autoregressive panel models with random effects. The estimation is performed by maximising the concentrated profile likelihood function by means of a non-linear optimisation algorithm. Eventually, Chang et al., (2010) and Hsu et al., (2012) additionally reduced the model’s complexity by combining covariance tapering and PMLE for spatial and spatiotemporal settings, respectively.
4.2 Regularisation for the spatial dependence structure
@@
Alternatively, the spatial interactions can be modelled in the random effects term in the same manner, while the errors remain independent across space. Generally, there is a relation between geostatistical and spatial autoregression models, and both approaches are equivalent under certain conditions (Ver Hoef et al., 2018a, , Theorem 1).
@@
Instead of the linear relation $\rho\mathbf{W}$, the spatial interactions can be modelled using a series of different weight structures, e.g., $\sum_{i=1}^{k}\rho_{i}\mathbf{W}_{i}$ with $k$ different weight matrices $\mathbf{W}_{i}$. For instance, each weight matrix could only contain the weights for certain directions (northward, north-eastward, eastward dependence, etc.) to reveal directional processes (Merk and Otto,, 2021). Moreover, penalised estimation procedures can be used to select the true weight matrix $\mathbf{W}$ from a series of alternative weights $\mathbf{W}_{1},\ldots,\mathbf{W}_{k}$, as for the boosting procedure proposed by Kostov, (2010, 2013) or the LASSO least-squares procedure proposed by Lam and Souza, (2020). Reyes et al., (2012) applied a spatiotemporal LASSO procedure to select weight matrices from a set of candidates with increasing spatial lag order, simultaneously with the temporal lags and spatiotemporal weight matrices. In other words, they constructed a 2-dimensional grid of the temporal and spatial lag orders and selected the relevant spatiotemporal interactions. In the context of traffic analysis, Haworth and Cheng, (2014) applied a graphical LASSO procedure to select local neighbourhood structures analogously. Together with the autoregressive coefficients, Reyes et al., (2012) also penalised the regressive parameters using a second penalty term. Similarly, Liu et al., (2018) suggested a LASSO procedure for selecting the regressors in the mean equation while allowing for spatial autoregressive structure in the model. Since ordinary least-squares procedures are inconsistent in the presence of spatial autoregressive dependence, they proposed a penalised quasi-maximum likelihood approach incorporating a LASSO penalty with a zero shrinkage target for the regression coefficients.
@@ 5 Cross validation under spatiotemporal dependence
When a strictly positive $\lambda$ is used, the estimated coefficients $\hat{\beta}_{\lambda}$ are biased but are more efficient, i.e., their variability is smaller. This is known as the bias-variance trade-off (see, for instance, Section 2.9 of Hastie et al.,, 2009). Additionally, opting for a $\lambda$ value that is too small can result in overfitting, while selecting a value that is too large can lead to underfitting (see, e.g., Boonstra et al.,, 2015). The optimal penalty parameter is usually selected based on the model’s in-sample or out-of-sample predictive performance. Other ways to select the penalty parameter are likelihood-based methods, where $\lambda$ is interpreted as a variance component and the likelihood is maximised with respect to the couple $(\sigma_{\varepsilon}^{2},\lambda)^{\prime}$. Similarly, the penalty parameter can be simultaneously estimated with all parameters in a fully Bayesian approach, where typically vague half-Cauchy prior distribution is assumed for $\lambda$ (van Erp et al.,, 2019). Alternatively, Otto and Steinert, (2023) proposed to select the penalty parameter based on the distance between the sample and model spatial autocorrelation. Below, we will focus on the predominantly applied goodness-of-fit (GoF) criteria.
GoF criteria aim to maximise the model fit by tuning the penalty parameter $\lambda$. The model fit may be assessed either in-sample or, better, out-of-sample, typically using cross-validation in terms of the distance between the observed and predicted values. The distance is typically evaluated by Root-Mean-Squared-Error (RMSE), Mean Absolute Error (MAE), or based on information criteria, such as Akaike’s Information Criterion (Akaike,, 1973, AIC) and the Bayesian Information Criterion (Schwarz,, 1978, BIC). Most models have no closed-form solutions to determine the optimal value of $\lambda$, necessitating grid-search algorithms. Arlot and Celisse, (2010) offers a thorough review of the significance of cross-validation (CV) in regression model selection while providing guidelines to choose the suitable cross-validation procedure according to the specificities of the data, including some brief remarks on dependent data and the problem at hand (e.g., model identification or model selection).
@@
    Mutual independence of test errors;
@@
    Independence between the training and test sets.
Regarding the first point, if a model were trained on peak points of seasonal time series and tested on the valley points, the prediction errors would be overestimated. Thus, while periodic partitioning should be avoided, a random partitioning strategy should always be preferred. For illustration of the other two points, consider overlapping folds or partitions. In such a situation, training and test sets would be mutually correlated, as well as the resulting training and test errors. As a direct consequence, the sample variance estimator would remarkably underestimate the actual variance of test errors. Independence among the sets can be assured by leaving a certain distance between training and test samples. Thus, for a given test set, all other correlated samples have to be removed from the training set to avoid overfitting It means, for example, that when considering spatial or temporal data, the nearby measurements (locations in space or time points) should be removed from the test set when validating a point.
@@
Depending on the data structure (e.g., cross-sectional, clustered, spatial, temporal data), as well as on modelling purposes and data-specific features (e.g., seasonality or non-Gaussianity), the cross-validation schemes to be adopted and may vary significantly (Arlot and Celisse,, 2010; Hewamalage et al.,, 2023). In the case of independent data, classical CV schemes, such as random $k$-fold, stratified $k$-fold (Zeng and Martinez,, 2000; Ludwig et al.,, 2016) or Generalized CV (Boonstra et al.,, 2015) can be used. For instance, considering ridge regularisation, the GCV estimator can be used to efficiently estimate the penalty parameter $\lambda$ even when the number of observations (thus, the degrees of freedom) is small or the number of parameters to be estimated exceeds the number of observations (Golub et al.,, 1979). For spatially and temporally dependent data, standard random $k$-fold cross-validation procedures should not be applied. As shown by Schratz et al., (2019), while performances are overestimated when no spatial information is included in the CV step, hyperparameter tuning of machine learning models appears to be less sensitive to the spatial structure, leading to similar results of non-spatial and spatial CV schemes.
@@
The problems arising from the occurrence of spatiotemporal autocorrelation in model assessment are manifold. First, using random sampling in cross-validation results in test observations being collected from areas that are spatially close to the training observations (Schratz et al.,, 2019). Consequently, the evaluation of prediction performance tends to be overly optimistic because the training and test datasets become correlated, largely due to the neglected underlying correlation structure, whether across space or time (Brenning,, 2012; Meyer et al.,, 2019; Meyer and Pebesma,, 2021; Ploton et al.,, 2020; Lezama Valdes et al.,, 2021). That is, if one leaves out a set of observations and estimates a spatiotemporal model with the remaining observations, information from the observations used for model estimation is used to predict the left-out observations (via the spatiotemporal interactions). Second, when the spatiotemporal structure is neglected, the estimated residuals will not be mutually independent, which is a critical assumption in many statistical models. As a result, it is often advisable to exclude complete data blocks across time and/or space (Roberts et al.,, 2017).
5.1 Cross-validation in time